# Automating Continual Learning

### v1

@@ 6 Conclusion
Our novel Automated Continual Learning (ACL) formulates continual learning as sequence learning across long time lags, and trains sequence-processing self-referential neural networks (SRNNs) to learn their own in-context continual (meta-)learning algorithms. ACL encodes classic desiderata for continual learning (such as knowledge preservation, forward and backward transfer, etc.) into the objective function of the meta-learner. ACL uses gradient descent to deal with classic challenges of CL, to automatically discover CL algorithms with good behavior. Once trained, our SRNNs autonomously run their own continual learning algorithms without requiring any human intervention. Our experiments demonstrate the effectiveness and promise of the proposed approach.
@@ A.1 Continual and meta-learning Terminologies
Here we review the following classic terminologies of continual learning and meta-learning used throughout this paper.
@@
“Domain-incremental learning (DIL)” and “class-incremental learning (CIL)” are two classic settings in continual learning (Van de Ven & Tolias, 2018a; b; Hsu et al., 2018). They differ as follows. Let $M$ and $N$ denote positive integers. Consider continual learning of $M$ tasks where each task is an $N$-way classification. In the DIL case, a model has an $N$-way output classification layer, i.e., the class ‘0’ of the first task shares the same weights as the class ‘0’ of the second task, and so on. In the CIL case, a model’s output dimension is $N*M$; the class indices of different tasks are not shared, neither are the corresponding weights in the output layer. In our experiments, all CIL models have the $(N*M)$-way output from the first task (instead of progressively increasing the output size). In this work, we skip the third variant called “task-incremental learning” which assumes that we have access to the task identity as an extra input, as it makes the CL problem almost trivial. CIL is typically reported to be the hardest setting among them.
@@ A.2 Datasets
We use torchmeta (Deleu et al., 2019) which provides common few-shot/meta learning settings for these datasets to sample and construct their meta-train/test datasets. The construction of “meta-training training” sequences for an $N$-way classification, using a dataset containing $C$ classes works as follows; for each sequence, we sample $N$ random but distinct classes out of $C$ ($N<C$). The resulting classes are re-labelled such that each class is assigned to one out of $N$ distinct random label index which is unique to the sequence. For each of these $N$ classes, we sample $K$ examples. We randomly order these $N*K$ examples to obtain a sequence. Each such a sequence “simulates” an unknown task the model has to learn.
@@ A.4 Evaluation Procedure
For the Split-MNIST experiment, we do 10 meta-testing runs to compute the mean and standard deviation as the baseline models are also trained for 10 runs in Hsu et al. (2018) (see other details in Appendix A.7.
A.5 Three-Task ACL
@@
$-\left(\log(p(y^{\mathcal{C}}_{\text{test}}|{\bm{x}}^{\mathcal{C}}_{\text{test}};{\bm{W}}_{\mathcal{A,B,C}}))+\log(p(y^{\mathcal{B}}_{\text{test}}|{\bm{x}}^{\mathcal{B}}_{\text{test}};{\bm{W}}_{\mathcal{A,B,C}}))+\log(p(y^{\mathcal{A}}_{\text{test}}|{\bm{x}}^{\mathcal{A}}_{\text{test}};{\bm{W}}_{\mathcal{A,B,C}}))\right)$ (5)
@@
This also naturally extends to the 5-task loss used in the Split-MNIST experiment (Table 3). As one can observe, the number of terms rapidly/quadratically increases with the number of tasks. Nevertheless, computing these loss terms isn’t immediately impractical because they essentially just require forwarding the network for one step, for many independent inputs/images. This can be heavily parallelized as a batch operation. While this can be a concern when scaling up more, a natural open research question is whether we really need all these terms in the case we have many more tasks. Our experiments effectively show that using more terms in the ACL loss improves performance. That said, ideally, we want these models to ‘systematically generalize’ to more tasks even when they are trained with only a handful of them (Fodor & Pylyshyn, 1988). This as an interesting research question on generalization to be studied in a future work.
A.6 Auxiliary 1-shot Learning Objective
A.7 Details of the Split-MNIST experiment
Split-MNIST is obtained by transforming the classic 10-class single-task MNIST dataset into a sequence of 5 tasks by partitioning the 10 classes into 5 groups/pairs of two classes each, in a fixed order from 0 to 9 (i.e., grouping 0/1, 2/3, 4/5, 6/7, and 8/9). For the difference between domain/class-incremental settings, we refer to Appendix A.1.
For the 2-task class incremental setting, we use Hsu et al. (2018)’s code to train the correspond models (the number for LwF is currently missing as it is not implemented in their code base; we will add the corresponding/missing entry in Table 3 for the final version of this paper).
Finally we also evaluate a MAML-based meta-learning approach, OML (Javed & White, 2019). We note that as reported by Javed & White (2019) in their public code repository; after some critical bug fix, the performance of their OML matches that of Beaulieu et al. (2020) (which is a direct application of OML to another model architecture). Therefore, we focus on OML as our main meta-continual learning baseline. We take the out-of-the-box model (meta-trained for Omniglot, with a 1000-way output) made publicly available by (Javed & White, 2019). We evaluate the corresponding model in two ways. In the first, ‘out-of-the-box’ case, we take the meta-pre-trained model and only tune its meta-testing learning rate (which is done by Javed & White (2019) even for meta-testing in Omniglot). We find that this setting does not perform very well; in the other case (‘optimized # meta-testing iterations’), we additionally tune the number of meta-test training iterations. We’ve done a grid search of the meta-test learning rate in $3*\{1e^{-2},1e^{-3},1e^{-4},1e^{-5}\}$ and the number of meta-test training steps in $\{1,2,5,8,10\}$ using a meta-validation set based on an MNIST validation set (5 K held-out images from the training set); we found the learning rate of $3e^{-4}$ and $8$ steps to consistently perform the best in all our settings. We’ve also tried it ‘with’ and ‘without’ the standard mean/std normalization of the MNIST dataset; better performance was achieved without such normalization (which is in fact consistent as they do not normalize the Omniglot dataset for their meta-training/testing). Their performance on the 5-task class-incremental setting is somewhat surprising/disappointing (since genenralization from Omniglot to MNIST is typically straightforward, at least, in common non-continual few-shot learning settings; see, e.g., Munkhdalai & Yu (2017)). At the same time, to the best of our knowledge, OML-trained models have not been tested in such a condition in prior work; from what we observe, the publicly available out-of-the-box model might be overtuned for Omniglot/Mini-ImageNet or the frozen ‘representation network’ is not ideal for genenralization. We note that the sensitivity of these MAML-based methods (Javed & White, 2019; Beaulieu et al., 2020) w.r.t. meta-test hyper-parameters has been also noted by Banayeeanzade et al. (2021); these are characteristics of hand-crafted learning algorithms that we want to avoid with learned learning algorithms.
Unlike any other methods above, our out-of-the-box ACL models (trained on Omniglot and Mini-ImageNet) do not require any tuning at meta-test time. Nevertheless, we’ve checked the effect of the number of meta-test training examples (5 vs. 15; 15 is the number used in meta-training); we found the consistent number, i.e., 15, to work better than 5. For the version that is meta-finetuned using the 5-task ACL objective (using only the Omniglot dataset), we use 5 examples for both meta-train and meta-test training. To obtain a sequence of 5 tasks, we simply sample 5 tasks from Omniglot (in principle, we should make sure that different tasks in the same sequence have no class overlap; in practice, our current implementation simply randomly draws 5 independent tasks from Omniglot).
Appendix B Extra Experimental Results and Comments
B.1 Ablation Studies on the choice of meta-validation sets
Here we conduct ablation studies on the choice of meta-validation sets to select model checkpoints. In general, when dealing with out-of-domain generalization, the choice of validation procedures to select final model checkpoints plays a crucial role in the evaluation of the corresponding method (Csordás et al., 2021; Irie et al., 2021b). The out-of-the-box models are chosen based on the average meta-validation performance on the validation set corresponding to the few-shot learning datasets used in meta-training: Omniglot and mini-ImageNet (or Omniglot, mini-ImageNet, and FC100 in the case of 3-task ACL), independently of any potential meta-test datasets. In contrast, in the meta-finetuning process of Table 3, we selected our model checkpoint by meta-validation on the MNIST validation dataset (we held out 5 K images from the training set). Here we evaluate ACL models meta-finetuned for the “5-task domain-incremental binary classification” on three Split-’X’ tasks where ’X’ is MNIST, FashionMNIST (FMNIST) or CIFAR-10 for various choices of meta-validation sets (in each case we hold out 5 K images from the corresponding training set). In addition, we also evaluate the effect of meta-finetuning datasets (Omniglot only v. Omniglot and mini-ImageNet). Table 6 shows the results (we use 15 meta-training and meta-testing examples except for the Omniglot-finedtuned/MNIST-validated model from Table 3 which happens to be configured with 5 examples; this will be fixed in the final version). Effectively, meta-validation on the matching validation set is useful. Also, meta-finetuning only on Omniglot is beneficial for the performance on MNIST when meta-validated on MNIST or FMNIST. However, importantly, we emphasize that our ultimate goal is not to obtain a model that is specifically tuned for certain datasets; we aim at building models that generally work well across a wide range of tasks (ideally on any tasks); in fact, several existing works in the few-shot learning literature evaluate their methods in such settings (see, e.g., Requeima et al. (2019); Bronskill et al. (2020); Triantafillou et al. (2020)). This also goes hand-in-hand with scaling up ACL (our current model is tiny; see hyper-parameters in Table 5; the vision component is also a shallow ‘Conv-4’ net) and various other considerations on self-improving continual learners (see, e.g., Schmidhuber (2018)), such as automated curriculum learning (Graves et al., 2017).
@@ B.1 Ablation Studies on the choice of meta-validation sets
Table 6: Impact of the choice of meta-validation datasets. Classification accuracies (%) on three datasets: Split-CIFAR-10, Split-Fashion MNIST (Split-FMNIST), and Split-MNIST in the domain-incremental setting. “OOB” denotes “out-of-the-box”. “mImageNet” here refers to mini-ImageNet.
@@
B.2 A Comment on Meta-Generalization

### v2

@@ 6 Conclusion
Our Automated Continual Learning (ACL) formulates continual learning as sequence learning across long time lags, and trains sequence-processing self-referential neural networks (SRNNs) to learn their own in-context continual (meta-)learning algorithms. ACL encodes classic desiderata for continual learning (such as knowledge preservation, forward and backward transfer, etc.) into the objective function of the meta-learner. ACL uses gradient descent to deal with classic challenges of CL, to automatically discover CL algorithms with good behavior. Once trained, our SRNNs autonomously run their own CL algorithms without requiring any human intervention. Our experiments demonstrate the effectiveness and promise of the proposed approach, including classic CL benchmark settings where existing hand-crafted algorithms fail.
@@ A.1 Continual and Meta-learning Terminologies
We review the following classic terminologies of continual learning and meta-learning used throughout this paper.
@@
“Domain-incremental learning (DIL)” and “class-incremental learning (CIL)” are two classic settings in continual learning (Van de Ven & Tolias, 2018a, b; Hsu et al., 2018). They differ as follows. Let $M$ and $N$ denote positive integers. Consider continual learning of $M$ tasks where each task is an $N$-way classification. In the DIL case, a model has an $N$-way output classification layer, i.e., the class ‘0’ of the first task shares the same weights as the class ‘0’ of the second task, and so on. In the CIL case, a model’s output dimension is $N*M$; the class indices of different tasks are not shared, neither are the corresponding weights in the output layer. In our experiments, all CIL models have the $(N*M)$-way output from the first task (instead of progressively increasing the output size). In this work, we skip the third variant called “task-incremental learning” which assumes that we have access to the task identity as an extra input, as it makes the CL problem almost trivial. CIL is typically reported to be the hardest setting among them.
@@ A.2 Datasets
Meta-train/test sequence construction procedure. We use torchmeta (Deleu et al., 2019) which provides common few-shot/meta learning settings for these datasets to sample and construct their meta-train/test datasets. The construction of “meta-training training” sequences for an $N$-way classification, using a dataset containing $C$ classes works as follows; for each sequence, we sample $N$ random but distinct classes out of $C$ ($N<C$). The resulting classes are re-labelled such that each class is assigned to one out of $N$ distinct random label index which is unique to the sequence. For each of these $N$ classes, we sample $K$ examples. We randomly order these $N*K$ examples to obtain a sequence. Each such a sequence “simulates” an unknown task the model has to learn.
@@ A.4 Evaluation Procedure
For the Split-MNIST experiment, we do 10 meta-testing runs to compute the mean and standard deviation as the baseline models are also trained for 10 runs in Hsu et al. (2018) (see other details in Appendix A.7).
A.5 ACL Objectives with More Tasks
@@
$-\Bigl{(}\log(p(y^{\mathcal{C}}_{\text{test}}|{\bm{x}}^{\mathcal{C}}_{\text{test}};{\bm{W}}_{\mathcal{A,B,C}}))$ (5)
$+\log(p(y^{\mathcal{B}}_{\text{test}}|{\bm{x}}^{\mathcal{B}}_{\text{test}};{\bm{W}}_{\mathcal{A,B,C}}))+\log(p(y^{\mathcal{A}}_{\text{test}}|{\bm{x}}^{\mathcal{A}}_{\text{test}};{\bm{W}}_{\mathcal{A,B,C}}))\Bigr{)}$
@@
This also naturally extends to the 5-task loss used in the Split-MNIST experiment (Table 3). As one can observe, the number of terms rapidly/quadratically increases with the number of tasks. Nevertheless, computing these loss terms isn’t immediately impractical because they essentially just require forwarding the network for one step, for many independent inputs/images. This can be heavily parallelized as a batch operation. While this can be a concern when scaling up more, a natural open research question is whether we really need all these terms in the case we have many more tasks. Ideally, we want these models to ‘systematically generalize’ to more tasks even when they are trained with only a handful of them (Fodor & Pylyshyn, 1988). This is an interesting research question on generalization to be studied in a future work.
@@
Table 6: Impact of the choice of meta-validation datasets. Classification accuracies (%) on three datasets: Split-CIFAR-10, Split-Fashion MNIST (Split-FMNIST), and Split-MNIST in the domain-incremental setting (we omit “Split-” in the second column). “OOB” denotes “out-of-the-box”. “mImageNet” here refers to mini-ImageNet.
@@
A.6 Auxiliary 1-shot Learning Objective
A.7 Details of the Split-MNIST experiment
Split-MNIST is obtained by transforming the classic 10-class single-task MNIST dataset into a sequence of 5 tasks by partitioning the 10 classes into 5 groups/pairs of two classes each, in a fixed order from 0 to 9 (i.e., grouping 0/1, 2/3, 4/5, 6/7, and 8/9). Regarding the difference between domain/class-incremental settings, we refer to Appendix A.1.
For the 2-task class incremental setting, we use Hsu et al. (2018)’s code to train the corresponding models (the number for LwF is currently missing as it is not implemented in the corresponding code base).
Finally we also evaluate two meta-CL baselines: Online-aware Meta-Learning (OML; Javed & White (2019)) and Generative Meta-Continual Learning (GeMCL; Banayeeanzade et al. (2021)). OML is a MAML-based meta-learning approach. We note that as reported by Javed & White (2019) in their public code repository; after some critical bug fix, the performance of their OML matches that of Beaulieu et al. (2020) (which is a direct application of OML to another model architecture). Therefore, we focus on OML as our main MAML-based baseline. We take the out-of-the-box model (meta-trained for Omniglot, with a 1000-way output) made publicly available by Javed & White (2019). We evaluate the corresponding model in two ways. In the first, ‘out-of-the-box’ case, we take the meta-pre-trained model and only tune its meta-testing learning rate (which is done by Javed & White (2019) even for meta-testing in Omniglot). We find that this setting does not perform very well; in the other case (‘optimized # meta-testing iterations’), we additionally tune the number of meta-test training iterations. We’ve done a grid search of the meta-test learning rate in $3*\{1e^{-2},1e^{-3},1e^{-4},1e^{-5}\}$ and the number of meta-test training steps in $\{1,2,5,8,10\}$ using a meta-validation set based on an MNIST validation set (5 K held-out images from the training set); we found the learning rate of $3e^{-4}$ and $8$ steps to consistently perform the best in all our settings. We’ve also tried it ‘with’ and ‘without’ the standard mean/std normalization of the MNIST dataset; better performance was achieved without such normalization (which is in fact consistent as they do not normalize the Omniglot dataset for their meta-training/testing). Their performance on the 5-task class-incremental setting is somewhat surprising/disappointing (since genenralization from Omniglot to MNIST is typically straightforward, at least, in common non-continual few-shot learning settings; see, e.g., Munkhdalai & Yu (2017)). At the same time, to the best of our knowledge, OML-trained models have not been tested in such a condition in prior work; from what we observe, the publicly available out-of-the-box model might be overtuned for Omniglot/Mini-ImageNet or the frozen ‘representation network’ is not ideal for genenralization. We note that the sensitivity of these MAML-based methods (Javed & White, 2019; Beaulieu et al., 2020) w.r.t. meta-test hyper-parameters has been also noted by Banayeeanzade et al. (2021); these are characteristics of hand-crafted learning algorithms that we want to avoid with learned learning algorithms.
We use code and a pre-trained model (trained on Omniglot) made public by Banayeeanzade et al. (2021) for the GeMCL baseline (see also Table 7); like our method, GeMCL also do not require any special tuning at test-time.
Our out-of-the-box ACL models (trained on Omniglot and Mini-ImageNet) do not require any tuning at meta-test time. Nevertheless, we’ve checked the effect of the number of meta-test training examples (5 vs. 15; 15 is the number used in meta-training); we found the consistent number, i.e., 15, to work better than 5. For the version that is meta-finetuned using the 5-task ACL objective (using only the Omniglot dataset), we use 5 or 15 examples for both meta-train and meta-test training (see an ablation study in Table 7). To obtain a sequence of 5 tasks, we simply sample 5 tasks from Omniglot (in principle, we should make sure that different tasks in the same sequence have no class overlap; in practice, our current implementation simply randomly draws 5 independent tasks from Omniglot).
Table 7: Impact of the number of in-context examples. Classification accuracies (%) on Split-MNIST in the 2-task and 5-task class-incremental learning (CIL) settings and the 5-task domain-incremental learning (DIL) setting. For ACL models, we use the same number of examples for meta-validation as for meta-training. According to Banayeeanzade et al. (2021), GeMCL is meta-trained with the 5-shot setting but meta-validated in the 15-shot setting.
Number of Examples DIL CIL 2-task CIL 5-task
Meta-Train/Valid Meta-Test GeMCL ACL GeMCL ACL GeMCL ACL
5 5 - 84.1 $\pm$ 1.2 - 93.4 $\pm$ 1.2 - 74.6 $\pm$ 2.3
15 - 83.8 $\pm$ 2.8 - 94.3 $\pm$ 1.9 - 65.5 $\pm$ 4.0
15 5 62.2 $\pm$ 5.2 83.9 $\pm$ 1.0 87.3 $\pm$ 2.5 93.6 $\pm$ 1.7 71.7 $\pm$ 2.5 76.7 $\pm$ 3.6
15 63.8 $\pm$ 3.8 84.5 $\pm$ 1.6 91.2 $\pm$ 2.8 96.0 $\pm$ 1.0 79.0 $\pm$ 2.1 84.3 $\pm$ 1.2
Table 8: Meta-testing on sequences that are longer than those from meta-training. Classification accuracies (%) on 5-task Split-FMNIST and 5-task Split-MNIST in the domain-incremental settings. The model is the one finetuned with 5-task ACL loss using Omniglot as the meta-finetuning set and FMNIST as the meta-validation set (i.e., the numbers in the top part of the table are taken from Table 6). In the first column, “Split-FMNIST, Split-MNIST” indicates continual learning of 5 Split-FMNIST tasks followed by 5 tasks of Split-MNIST (and “Split-MNIST, Split-FMNIST” is the opposite order). Performance is measured at the end of the entire sequence.
Meta-Test Test Tasks
Meta-Test Training Task Sequence # Tasks Split-FMNIST Split-MNIST
Split-FMNIST 5 90.4 $\pm$ 0.5 -
Split-MNIST 5 - 81.6 $\pm$ 1.3
Split-FMNIST, Split-MNIST 10 79.3 $\pm$ 2.7 74.3 $\pm$ 0.9
Split-MNIST, Split-FMNIST 10 78.1 $\pm$ 3.1 78.5 $\pm$ 1.7
Appendix B Extra Experimental Results
B.1 Ablation Studies on the Meta-validation Dataset
Here we conduct ablation studies on the choice of meta-validation sets to select model checkpoints. In general, when dealing with out-of-domain generalization, the choice of validation procedures to select final model checkpoints plays a crucial role in the evaluation of the corresponding method (Csordás et al., 2021; Irie et al., 2021b). The out-of-the-box models are chosen based on the average meta-validation performance on the validation set corresponding to the few-shot learning datasets used in meta-training: Omniglot and mini-ImageNet (or Omniglot, mini-ImageNet, and FC100 in the case of 3-task ACL), independently of any potential meta-test datasets. In contrast, in the meta-finetuning process of Table 3, we selected our model checkpoint by meta-validation on the MNIST validation dataset (we held out 5 K images from the training set). Here we evaluate ACL models meta-finetuned for the “5-task domain-incremental binary classification” on three Split-‘X’ tasks where ‘X’ is MNIST, FashionMNIST (FMNIST) or CIFAR-10 for various choices of meta-validation sets (in each case we hold out 5 K images from the corresponding training set). In addition, we also evaluate the effect of meta-finetuning datasets (Omniglot only v. Omniglot and mini-ImageNet). Table 6 shows the results (we use 15 meta-training and meta-testing examples except for the Omniglot-finedtuned/MNIST-validated model from Table 3 which happens to be configured with 5 examples; this will be fixed in the final version). Effectively, meta-validation on the matching validation set is useful. Also, meta-finetuning only on Omniglot is beneficial for the performance on MNIST when meta-validated on MNIST or FMNIST. However, importantly, we emphasize that our ultimate goal is not to obtain a model that is specifically tuned for certain datasets; we aim at building models that generally work well across a wide range of tasks (ideally on any tasks); in fact, several existing works in the few-shot learning literature evaluate their methods in such settings (see, e.g., Requeima et al. (2019); Bronskill et al. (2020); Triantafillou et al. (2020)). This also goes hand-in-hand with scaling up ACL (our current model is tiny; see hyper-parameters in Table 5; the vision component is also a shallow ‘Conv-4’ net) and various other considerations on self-improving continual learners (see, e.g., Schmidhuber (2018)), such as automated curriculum learning (Graves et al., 2017).
B.2 Performance on Split-Omniglot
Here we report the performance of the models used in the Split-MNIST experiment (Sec. 4.3) on “in-domain” 5-task 2-way Split-Omniglot. Table 9 shows the result. Performance is very similar between our ACL and the baseline GeMCL on this task in the class incremental setting, unlike on Split-MNIST (Table 3) where we observe a larger performance gap between these same models. Here we also include the “domain incremental” setting for the sake of completeness but note that GeMCL is not originally trained for this setting.
Table 9: Classification acuracies (%) on 5-task 2-way Split-Omniglot. Mean/std is computed over 10 meta-test runs.
Method Domain Incremental Class Incremental
GeMCL 64.6 $\pm$ 9.2 97.4 $\pm$ 2.7
ACL 92.3 $\pm$ 0.4 96.8 $\pm$ 0.8
B.3 Effect of Number of In-Context Examples
Table 7 shows an ablation study on the number of examples used for meta-training and meta-testing on the Split-MNIST task. We observe that for an ACL model trained only with 5 examples during meta-training, more examples (15 examples) provided during meta-testing is not beneficial. In fact, they even largely hurt in certain cases (see the last column); this is one form of “length generalization” problem. When the number of meta-training examples is consistent with the one used during meta-testing, the 15-example case consistently outperforms the 5-example one.
B.4 Further Discussion on Limitations
Here we provide further discussion and experimental results on the limitations of our approach as a learned algorithm.
Domain generalization.
As a data-driven learned algorithm, the domain generalization capability is a typical limitation as it depends on the meta-trained data. Certain results we presented above are representative of this limitation. In particular, in Table 6, the model meta-trained/finetuned on Omniglot using Split-MNIST as meta-validation set do not perform well on Split-CIFAR10. While meta-training and meta-validating on a larger/diverse set of datasets may be an immediate remedy to obtain more robust ACL models, we note that since ACL is also a “continual meta-learning” algorithm (Sec. 5), an ideal ACL model should also continually incorporate and learn from more data during potentially lifelong meta-testing; we leave such an investigation for future work.
Length generalization.
We already qualitatively observed the limited length generalization capability in Table 4 (meta-trained with up to 3 tasks and meta-tested with up to 4 tasks). Here we provide one more experiment evaluating ACL models meta-trained for 5 tasks on a concatenation of two 5-task Split-MNIST and Split-FMNIST tasks (resulting in 10 tasks). Table 8 shows the results. Again, while the model does not completely break, increasing the number of tasks to 10 rapidly degrades the performance compared to the 5-task setting the model is meta-trained for. Similarly, its performance on the Split-Omniglot domain incremental setting (Sec. B.2) degrades with increased numbers of tasks: accuracies for 5, 10 and 20 tasks are $92.3\%\pm 0.4$, $82.0\%\pm 0.4$ and $67.6\%\pm 1.1$ respectively. As noted in Sec. 5, this is a general limitation of sequence processing neural networks, and there is a potential remedy for this limitation (meta-training on more tasks and “context carry-over”) which we leave for future work.
B.5 A Comment on Meta-Generalization